1.  Ben has determined that the length of time it takes him to commute from his residence to school is normally distributed with a mean of 45 minutes and a standard deviation of 5 minutes. How much time must he allow if he wishes to be in his first class on time 90% of the time?



1
Expert's answer
2021-09-21T00:29:43-0400

Let "X=" the length of time: "X\sim N(\mu, \sigma^2)."

Given "\mu=45\ min, \sigma=5\ min."


"P(X<x)=P(Z<\dfrac{x-\mu}{\sigma})=P(Z<\dfrac{x-45}{5})=0.9"

"=P(Z<\dfrac{x-45}{5})=0.9"

"\dfrac{x-45}{5}\approx1.2816"

"x=51.4\ min"